OpenStudy (anonymous):

Please check my work. write a^-3b^-1c^2d^4 with positive exponents. would it be: 1/a^3, 1/b^1 c^2 d^4??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yay!!! tytyty

OpenStudy (anonymous):

In solving an equation for x, you distribute a –4 on one side of the equation. What needs to be done to the other side of the equation? A. multiply by 4 B. add –4 C. subtract 4 D. nothing

OpenStudy (mertsj):

So you would write it like this: \[\frac{c^2d^4}{a^3b}\]

OpenStudy (anonymous):

oh boy then i don't get it :(

OpenStudy (anonymous):

if it is originally in a line, why would it become a fraction?

OpenStudy (mertsj):

because a^-n=1/a^n

OpenStudy (anonymous):

ok that does make sense now,,,

OpenStudy (anonymous):

you are a best!!!

OpenStudy (mertsj):

\[a ^{-3}=\frac{1}{a^3}``b ^{-1}=\frac{1}{b}\]

OpenStudy (anonymous):

yes!

OpenStudy (mertsj):